OpenStudy (anonymous):

How do I get f(x) when I have f'(x)=x(1/2)?

OpenStudy (anonymous):

integrate

OpenStudy (zepp):

^

OpenStudy (anonymous):

\[\int\frac{x}{2}dx\]

OpenStudy (anonymous):

I'm confused how to integrate it though.

OpenStudy (anonymous):

Power Rule for integration \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (klimenkov):

Dont forget to except \(n=-1\) from your Power Rule.

OpenStudy (anonymous):

Okay thankies guys :)